# Ellipsoidal Density-Equalizing Map for Genus-0 Closed Surfaces

### v1

@@ 5 Application to genus-0 surface remeshing
In Figs. 11, 12, 13, and 14, we consider various genus-0 surface models and compare the remeshing results obtained by different parameterization methods for genus-0 closed surfaces, including spherical conformal mapping (SCM) [9], spherical density-equalizing mapping (SDEM) [43], ellipsoidal conformal mapping (ECM) [30], our ellipsoidal density-equalizing mapping (EDEM) method, and our ellipsoidal density-equalizing quasi-conformal mapping (EDEQ) method. For each genus-0 surface model $\mathcal{M}$ in $\mathbb{R}^{3}$ and each parameterization method, denote the parameterization mapping by $f:\mathcal{M}\rightarrow\mathcal{T}$, where the parameter domain $\mathcal{T}$ is either the unit sphere (for SCM and SDEM) or an ellipsoid (for ECM, EDEM and EDEQ). Then we construct a regular triangular mesh $\mathcal{T}_{\Delta}$ on $\mathcal{T}$ using the DistMesh method [55]. Finally, using the inverse mappings $f^{-1}:\mathcal{T}\to\mathcal{M}$, the regular mesh $\mathcal{T}_{\Delta}$ can be mapped back onto $\mathcal{M}$, yielding a remeshed surface $f^{-1}(\mathcal{T}_{\Delta})$. For a fair comparison, we enforce the number of vertices of the triangular mesh generated by DistMesh (and hence the number of vertices of the final remeshing result) for each method to be approximately 8500 (with variation <±$<\pm 1\%$).
From the remeshing results in Figs. 11, 12, 13, and 14, it can be observed that the two conformal approaches (SCM and ECM) lead to significantly non-uniform remeshing results. Moreover, the shape features near the sharp regions are lost in the SCM and ECM results. This can be explained by the fact that the conformal mappings preserve angles but may yield large area distortions. By contrast, the remeshing results obtained by the density-equalizing approaches (SDEM, EDEM, and EDEQ) are much more uniform and have a higher mesh quality. The features located near the sharp regions are well-preserved.
@@
[…]
@@
Table 7 records the values of $\delta_{\text{size}}$ and $\delta_{\text{shape}}$ for all remeshing results shown in Figs. 11, 12, 13, and 14. For the two conformal approaches (SCM and ECM), it can be observed that the values of $\delta_{\text{shape}}$ are small but the values of $\delta_{\text{size}}$ are extremely large, which means the remeshing results obtained by those two methods are highly non-uniform in size. By contrast, the values of $\delta_{\text{size}}$ for the three density-equalizing approaches (SDEM, EDEM, and EDEQ) are small. This demonstrates the advantage of utilizing density-equalizing maps for surface remeshing. Furthermore, by considering both $\delta_{\text{size}}$ and $\delta_{\text{shape}}$ for the three density-equalizing methods, it can be observed that the EDEQ method gives the best overall remeshing performance among the three methods, which also matches our observations in Fig. 15.
@@ 6 Discussion
In this work, we have proposed a novel method for computing bijective ellipsoidal density-equalizing maps (EDEM) for genus-0 closed surfaces. Using this approach, we can efficiently obtain parameterizations with different desired mapping effects, including ellipsoidal area-preserving parameterizations and ellipsoidal parameterizations with controlled area change. Also, by considering a combined energy involving both a density-equalizing term and a quasi-conformal term, we can achieve ellipsoidal density-equalizing quasi-conformal maps (EDEQ) balancing the angle and area distortions. Experimental results on various ellipsoidal surfaces and genus-0 closed surface models have demonstrated the effectiveness of our proposed methods. Our proposed methods can be easily applied to surface remeshing of genus-0 closed surfaces, with the remeshing quality significantly improved when compared to other prior approaches.
We remark that while our proposed methods have considered both the density-equalizing property and the quasi-conformality, landmark-matching constraints have not been incorporated into our formulations. In the future, we plan to further extend our methods to facilitate landmark-matching surface parameterization and surface registration. Our models may also be combined with curvature-based methods to achieve curvature-based surface registrations with controlled area change for surfaces with extreme geometries.

### v2

@@ 5 Application to genus-0 surface remeshing
In Fig. 11, Fig. 12, Fig. 13, and Fig. 14, we consider various genus-0 surface models and compare the remeshing results obtained by different parameterization methods for genus-0 closed surfaces, including spherical conformal mapping (SCM) [9], spherical density-equalizing mapping (SDEM) [44], ellipsoidal conformal mapping (ECM) [30], our ellipsoidal density-equalizing mapping (EDEM) method, and our ellipsoidal density-equalizing quasi-conformal mapping (EDEQ) method. For each genus-0 surface model $\mathcal{M}$ in $\mathbb{R}^{3}$ and each parameterization method, denote the parameterization mapping by $f:\mathcal{M}\rightarrow\mathcal{T}$, where the parameter domain $\mathcal{T}$ is either the unit sphere (for SCM and SDEM) or an ellipsoid (for ECM, EDEM and EDEQ). Then we construct a regular triangular mesh $\mathcal{T}_{\Delta}$ on $\mathcal{T}$ using the DistMesh method [56]. Finally, using the inverse mappings $f^{-1}:\mathcal{T}\to\mathcal{M}$, the regular mesh $\mathcal{T}_{\Delta}$ can be mapped back onto $\mathcal{M}$, yielding a remeshed surface $f^{-1}(\mathcal{T}_{\Delta})$. For a fair comparison, we enforce the number of vertices of the triangular mesh generated by DistMesh (and hence the number of vertices of the final remeshing result) for each method to be approximately 8500 (with variation $\pm 1\%$).
From the remeshing results in Fig. 11, Fig. 12, Fig. 13, and Fig. 14, it can be observed that the two conformal approaches (SCM and ECM) lead to significantly non-uniform remeshing results. Moreover, the shape features near the sharp regions are lost in the SCM and ECM results. This can be explained by the fact that the conformal mappings preserve angles but may yield large area distortions. By contrast, the remeshing results obtained by the density-equalizing approaches (SDEM, EDEM, and EDEQ) are much more uniform and have a higher mesh quality. The features located near the sharp regions are well-preserved.
@@
$\delta_{\text{shape}}=\underset{i}{\text{mean}}(R_{i}).$ (45)
@@
Table 7 records the values of $\delta_{\text{size}}$ and $\delta_{\text{shape}}$ for all remeshing results shown in Fig. 11, Fig. 12, Fig. 13, and Fig. 14. For the two conformal approaches (SCM and ECM), it can be observed that the values of $\delta_{\text{shape}}$ are small but the values of $\delta_{\text{size}}$ are large, which means the remeshing results obtained by those two methods are highly non-uniform in size. By contrast, the values of $\delta_{\text{size}}$ for the three density-equalizing approaches (SDEM, EDEM, and EDEQ) are small. This demonstrates the advantage of utilizing density-equalizing maps for surface remeshing. Furthermore, by considering both $\delta_{\text{size}}$ and $\delta_{\text{shape}}$ for the three density-equalizing methods, it can be observed that the EDEQ method gives the best overall remeshing performance among the three methods, which also matches our observations in Fig. 15.
@@ 6 Discussion
In this work, we have proposed a novel method for computing bijective ellipsoidal density-equalizing maps (EDEM) for genus-0 closed surfaces. Using this approach, we can efficiently obtain parameterizations with different desired mapping effects, including ellipsoidal area-preserving parameterizations and ellipsoidal parameterizations with controlled area change. Also, by considering a combined energy involving both a density-equalizing term and a quasi-conformal term, we can achieve ellipsoidal density-equalizing quasi-conformal maps (EDEQ) balancing the angle and area distortions. Experimental results on various ellipsoidal surfaces and genus-0 closed surface models have demonstrated the effectiveness of our proposed methods.
Our proposed methods have broad applicability across various scientific and engineering domains. As shown in the previous section, our methods can be applied to surface remeshing for genus-0 closed surfaces with significantly improved mesh quality. Besides, the proposed methods can be applied to texture mapping in the field of computer graphics. In particular, since the proposed methods can reduce the area and angle distortions of the parameterization of genus-0 closed surfaces, high-quality texture mapping results can be effectively obtained. Also, as demonstrated by our experiments on the Hippocampus model, our proposed methods are well-suited for the analysis of complex anatomical structures in medical imaging. Moreover, analogous to the prior works on density-equalizing map projection, our EDEM and EDEQ methods can be naturally applied to cartogram creation, data visualization, and geophysical surface modeling.
We remark that while our proposed methods have considered both the density-equalizing property and the quasi-conformality, landmark-matching constraints have not been incorporated into our formulations. In the future, we plan to further extend our methods to facilitate landmark-matching surface parameterization and surface registration. Besides, note that some practical applications may involve curvature-based surface registrations, which use curvature to guide the registration process between surfaces. By combining curvature-based energies and our proposed density-based and quasi-conformal energies, it may be possible to extend our model to achieve curvature-based surface registrations with controlled area change for surfaces with extreme geometries. We also plan to perform more theoretical analyses on the convergence of our proposed algorithms in future work.